# Temporal Analysis of Drifting Hashtags in Textual Data Streams: A Graph-Based Application

### v1

@@ Abstract
Social media has played an important role since its emergence. People use the internet to express opinions about anything, making social media platforms a social sensor. Initially supported by Twitter, the hashtags are now in use on several social media platforms. Hashtags are helpful to tag, track, and group posts on similar topics. In this paper, we analyze hashtag drifts over time using concepts from graph analysis and textual data streams using the Girvan-Newman method to uncover hashtag communities in annual snapshots. More specifically, we analyzed the #mybodymychoice hashtag between 2018 and 2022. In addition, we offer insights about some hashtags found in the study. Furthermore, our approach can be useful for monitoring changes over time in opinions and sentiment patterns about an entity on social media. Even though the hashtag #mybodymychoice was initially coupled with women’s rights, abortion, and bodily autonomy, we observe that it suffered drifts during the studied period across topics such as drug legalization, vaccination, political protests, war, and civil rights. The year 2021 was the most significant drifting year, in which the communities detected suggest that #mybodymychoice significantly drifted to vaccination and Covid-19-related topics.
@@
[…]
[inst1]organization=Programa de Pós-graduação em Informática (PPGIa), Pontifícia Universidade Católica do Paraná (PUCPR),city=Curitiba, state=Paraná, country=Brazil
[…]
[inst2]organization=Instituto Federal de Santa Catarina (IFSC),city=Caçador, state=Santa Catarina, country=Brazil
[…]
[inst3]organization=Universidade Estadual de Ponta Grossa (UEPG),city=Ponta Grossa, state=Paraná, country=Brazil
@@ 1 Introduction
The popularization of the Internet made it possible for people to express themselves in several ways via posts on social media using text, audio, videos, and pictures. People can also expose opinions about any topic, e.g., politics, products, and sports, and engage in collective actions and social movements. Social media has also enabled social movements to take place worldwide. For instance, authors in Eltantawy and Wiest (2011) discuss the role of social media in the Egyptian Revolution in 2011. The authors in Harlow (2012) analyzed Guatemalan users’ posts on Facebook groups to evaluate whether social media helped the movement in defense of Rodrigo Rosenberg to go offline in 2010. In addition, the authors in Harlow (2012) named these digital movements ‘Activism 2.0’, also due to Web 2.0, which includes social media networks. Another example of such activism in social media is the ‘Arab Spring,’ also discussed in several papers Smidi and Shahin (2017); Wolfsfeld et al. (2013); Comunello and Anzera (2012).
@@
More recently, during the Covid-19 pandemic, several movements in favor and against vaccination and other aspects of the pandemic, e.g., lockdown, have emerged Sufi et al. (2022). In Menghini et al. (2022), the authors discussed the use of #mybodymychoice, a feminist slogan to defend women’s rights, used for opposing Covid-19 vaccination campaigns.
@@
Twitter is one of the most popular social media platforms. Around 500 million tweets are posted per day in 2022111https://www.dsayce.com/social-media/tweets-day/. Each post is known as tweet. Its ease of posting and sharing information made it popular also for famous people, such as artists, athletes, and politicians. As of 2017, Twitter limits the post to 280 characters in its free version. In addition, people can tag their posts by using hashtags, leading to a post categorization. Researchers that need to evaluate information from social sensors, e.g., public opinion, resort to data from social media Suprem et al. (2019).
@@
Inspired by the authors in Menghini et al. (2022), we also analyzed the Twitter stream regarding the hashtag #mybodymychoice, comprising the years between 2018 and 2022. The authors in Menghini et al. (2022) demonstrated how the hashtag mentioned before, originally used to defend the right to abortion and other bodily autonomy aspects, drifted over time, being also used in different topics. Our work technically analyzes the context evolution of the hashtag, originating from a textual data stream corresponding to #mybodymychoice without judging the validity of the related social causes or the drifts of this hashtag. In addition, different from Menghini et al. (2022), which split the stream using two events known beforehand, we leveraged the Girvan-Newman community detection algorithm Girvan and Newman (2002) between years to evaluate the communities found.
@@
This paper aims to analyze the drifts in the hashtag stream of #mybodymychoice using time slices corresponding to years between 2018 and 2022, inclusive. We applied the Girvan-Newman algorithm for community detection after each year was finished. We analyzed the most relevant, i.e., the biggest, communities by evaluating the hashtags and their contexts. We are interested in analyzing the #mybodymychoice’s distancing from its original use or context. In addition, we listed the five most frequent hashtags per year. We consider drifts to have occurred whenever the hashtag is used in different contexts than its original. Therefore, our research questions are (RQ1) “How the hashtag #mybodymychoice evolved from 2018 to 2022?”, and (RQ2) “What were the most significant hashtag drifts over time, considering the graph communities detected?”. RQ1 aims to uncover drifts and their importance regarding a conceptual similarity to the original use of #mybodymychoice. RQ2 aims to find which topics other than women’s rights, abortion, and bodily autonomy were frequent in the period analyzed.
@@
The contribution of this work is two-fold: (i) the analysis of the hashtag stream respecting the textual data stream concepts and using graph-based approaches, including community detection between years, and (ii) to the best of our knowledge, this is the first time the Girvan-Newman algorithm was applied to a hashtag graph generated in a streaming fashion. A by-product of this work is the dataset containing hashtag stream regarding the #mybodymychoice hashtag between 2018 and 2022. This work may be relevant to applications that monitor sentiments or opinions regarding an entity on social media to recognize changes in patterns and behaviors.
@@ 2.1 Hashtags
Hashtags are words starting with the hash character (#) used for social media post categorization. Hashtags, under the name of channel tags, were elaborated by Chris Messina in 2007 with the idea of creating groups on Twitter [^2] Available at: https://factoryjoe.com/2007/08/25/groups-for-twitter-or-a-proposal-for-twitter-tag-channels/, accessed on June 20th, 2023. Twitter adopted hashtags as a feature in 2009 Scott (2015), which proved helpful and, later, was also integrated into other social media platforms, such as Facebook [^3] Available at: https://about.fb.com/news/2013/06/public-conversations-on-facebook/. Accessed on June 20th, 2023. in 2013.
Twitter post categorization or grouping leads to virtual community creation. This enables discussions on several aspects, and sometimes, online social movements go offline Smidi and Shahin (2017). An example of a hashtag used in social discussion is #mybodymychoice. As pointed out in Menghini et al. (2022), the hashtag #mybodymychoice initially regarded women’s right to abortion and bodily autonomy. However, the authors in Menghini et al. (2022) found that between 2018 and 2021, this hashtag drifted to other topics, such as politics and civil rights.
This paper expands the analysis performed in Menghini et al. (2022) to 2022, considering hashtags co-occurrence and community detection via temporal graphs and textual streaming analysis. In our analysis, we intend to trace back the events that used particular hashtags to understand the hashtags’ occurrences. Therefore, this section presents textual data stream mining and its characteristics, graphs and their notation, and techniques for community detection in graphs.
@@ 2.2 Textual Data Stream Mining
Data streams are “an algorithmic abstraction that allows for real-time analytics” Gama et al. (2014). Data arrive sequentially (one by one or in small batches) and quickly in streaming settings. In addition, a data stream may be infinite and ephemeral, meaning that the data distribution may change over time, thus giving rise to a phenomenon named concept drift Bifet et al. (2018); Gama et al. (2014).
Therefore, to learn from data streams, i.e., data stream mining, some capabilities desired for machine learning methods include Bifet et al. (2018); Gama et al. (2014): (a) learning quickly from the data as it arrives, (b) discarding data within a short time after learning from it, (c) performing single-pass processes, and (d) detecting and adapting to concept drifts.
Textual data streams are specializations of traditional data streams Thuma et al. (2023), i.e., it comprises a sequence of texts arriving continuously at a fast pace. However, processing texts in textual data stream scenarios is more challenging given the natural language processing (NLP)-related activities regarding, for example, vocabulary maintenance, token standardization, and processing texts only once.
Our proposed approach focuses on processing textual data streams while accounting for the above-mentioned constraints. The exception regards graph community detection, detailed in Section 2.4, which is performed on a year-basis, and, thus, should be considered an offline step.
@@ 2.3 Graph
A graph $G$ is a collection of edges $E$ and vertices $V$. Each vertex can be linked to another vertex by an edge. Generally, vertices are visually represented as circles, and the edges as lines between vertices, as depicted in the example graph in Fig. 1. Edges are directed if the direction between vertices matters and are undirected otherwise. Edges may also be coupled with weights to denote additional information about the relationship between connected vertices.
@@
Graphs are relevant to denote entities and the relationships between them, being very useful in text analysis. For instance, to find relevant words in sentences, the vertices are the words, and the edges are their relationships Castillo et al. (2017). In our proposal, vertices denote hashtags, and edges denote the co-occurrence between the linked hashtags. The edges are not weighted.
@@ 2.4 Community Detection in Graphs
Community detection in graphs groups nodes (vertices) with frequent links into a community. Criminal community detection Sarvari et al. (2014), cattle trade movement Cardoso et al. (2018), and disease spread dynamics Salathé and Jones (2010) are applications that leverage community detection. Several community detection methods are available in the literature, such as the Louvain method Blondel et al. (2008), the Leiden algorithm Traag et al. (2019), and the Girvan-Newman algorithm Girvan and Newman (2002).
@@
Similar to clustering, community detection demands the graph partition into densely connected nodes and nodes from a certain community to be far from those of other communities Blondel et al. (2008). According to Traag et al. (2019), the Louvain method obtains weakly connected communities. To handle this characteristic, the authors in Traag et al. (2019) added one more phase to the Louvain algorithm to obtain better-connected communities.
@@
The Girvan-Newman algorithm Girvan and Newman (2002) for community detection on graphs performs an iterative edge elimination. To consider an edge for elimination, it must be part of the highest number of shortest paths between nodes, i.e., betweenness centrality. After removing these edges, the communities are obtained. More specifically, the algorithm works in four steps Girvan and Newman (2002): (a) for all edges in the graph, compute the betweenness; (b) remove the edge with the highest betweenness; (c) for all edges affected by the edge removal in step (b), re-compute the betweenness; (d) return to step (b) until there is no edge remaining. We resorted to the Girvan-Newman algorithm because it generates consistent communities Sheng et al. (2020). However, its trade-off lies in its complexity, which makes it not scalable for large networks Sheng et al. (2020).
@@
Considering that we are using a textual data stream limited by a window and that the communities are detected in an offline step, community quality is the most important. Therefore, the Girvan-Newman algorithm’s complexity is not a concern in this particular scenario.

### v2

@@ Abstract
Initially supported by Twitter, hashtags are now used on several social media platforms. Hashtags are helpful for tagging, tracking, and grouping posts on similar topics. In this paper, based on a hashtag stream regarding the hashtag #mybodymychoice, we analyze hashtag drifts over time using concepts from graph analysis and textual data streams using the Girvan-Newman method to uncover hashtag communities in annual snapshots between 2018 and 2022. In addition, we offer insights about some correlated hashtags found in the study. Our approach can be useful for monitoring changes over time in opinions and sentiment patterns about an entity on social media. Even though the hashtag #mybodymychoice was initially coupled with women’s rights, abortion, and bodily autonomy, we observe that it suffered drifts during the studied period across topics such as drug legalization, vaccination, political protests, war, and civil rights. The year 2021 was the most significant drifting year, in which the communities detected and their respective sizes suggest that #mybodymychoice had a significant drift to vaccination and Covid-19-related topics.
@@
journal: Expert Systems with Applicationsjournal: Expert Systems with Application
@@ 1 Introduction
The popularization of the Internet made it possible for people to express themselves in several ways via posts on social media using text, audio, videos, and pictures. People can also expose opinions about any topic, e.g., politics, products, and sports, and engage in collective actions and social movements. Social media has also enabled social movements to take place worldwide. For instance, authors in Eltantawy & Wiest [2011] discuss the role of social media in the Egyptian Revolution in 2011. The authors in Harlow [2012] analyzed Guatemalan users’ posts on Facebook groups to evaluate whether social media helped the movement in defense of Rodrigo Rosenberg to go offline in 2010. In addition, the authors in Harlow [2012] named these digital movements ‘Activism 2.0’, also due to Web 2.0, which includes social media networks. Another example of such activism in social media is the ‘Arab Spring,’ also discussed in several papers [Smidi & Shahin, 2017, Wolfsfeld et al., 2013, Comunello & Anzera, 2012].
@@
More recently, during the Covid-19 pandemic, several movements in favor and against vaccination and other aspects of the pandemic, e.g., lockdown, have emerged [Sufi et al., 2022]. In Menghini et al. [2022], the authors discussed the use of #mybodymychoice, a feminist slogan to defend women’s rights, used for opposing Covid-19 vaccination campaigns.
@@
Twitter is one of the most popular social media platforms. Around 500 million tweets are posted per day in 2023111https://www.omnicoreagency.com/twitter-statistics/. Each post is known as tweet. Its ease of posting and sharing information made it popular also for famous people, such as artists, athletes, and politicians. As of 2017, Twitter limits the post to 280 characters in its free version. In addition, people can tag their posts by using hashtags, leading to a post categorization. Researchers that need to evaluate information from social sensors, e.g., public opinion, resort to data from social media [Suprem et al., 2019].
@@
Inspired by the authors in Menghini et al. [2022], we also analyzed the Twitter stream regarding the hashtag #mybodymychoice, comprising the years between 2018 and 2022. The authors in Menghini et al. [2022] demonstrated how the hashtag mentioned before, originally used to defend the right to abortion and other bodily autonomy aspects, drifted over time, being also used in different topics. Our work technically analyzes the context evolution of the hashtag, originating from a textual data stream corresponding to #mybodymychoice without judging the validity of the related social causes or the drifts of this hashtag. In addition, different from Menghini et al. [2022], which split the stream using two events known beforehand, we leveraged the Girvan-Newman community detection algorithm [Girvan & Newman, 2002] between years to evaluate the communities found.
@@
This paper aims to analyze the drifts in the hashtag stream of #mybodymychoice using time slices corresponding to years between 2018 and 2022, inclusive. We applied the Girvan-Newman algorithm for community detection after each year was finished. We analyzed the most relevant, i.e., the biggest, communities by evaluating the hashtags and their contexts. We are interested in analyzing the #mybodymychoice’s distancing from its original use or context. In addition, we listed the five most frequent hashtags per year. We consider drifts to have occurred whenever the hashtag is used in contexts different from its original. Therefore, our research questions are (RQ1) “How did the hashtag #mybodymychoice evolve from 2018 to 2022?”, and (RQ2) “What were the most significant hashtag drifts over time, considering the graph communities detected?”. RQ1 aims to uncover drifts and their importance regarding a conceptual similarity to the original use of #mybodymychoice. RQ2 aims to find which topics other than women’s rights, abortion, and bodily autonomy were frequent in the period analyzed. Although detecting hashtag drifts has been addressed in the literature, most analyses indicate to be in batch, thus demanding all the data forehand. Therefore, our motivation is to provide an online method to analyze hashtag streams subject to hashtag drifts that could be directly coupled to a streaming source, such as a social media platform Application Programming Interface (API).
@@
The contribution of this work is three-fold: (i) the analysis of the hashtag stream respecting the textual data stream concepts and using graph-based approaches, including community detection between years, (ii) to the best of our knowledge, this is the first time the Girvan-Newman algorithm was applied to a hashtag graph generated in a streaming fashion, and (c) an online graph management method to maintain the hashtag graph concise and with the most relevant hashtags. A by-product of this work is the dataset containing hashtag stream regarding the #mybodymychoice hashtag between 2018 and 2022. This work may be relevant to applications that monitor sentiments or opinions regarding an entity on social media to recognize changes in patterns and behaviors. The innovation of this work lies in the incremental graph generation method, which allows: (a) the management of the graph conciseness while preserving important hashtags, (b) the direct coupling to a text streaming API for online analysis, and (c) analyses over time, whenever it is needed.
@@ 2.1 Hashtags
Hashtags are words starting with the hash character (#) used for social media post categorization. Hashtags, under the name of channel tags, were elaborated by Chris Messina in 2007 with the idea of creating groups on Twitter [^2] Available at: https://factoryjoe.com/2007/08/25/groups-for-twitter-or-a-proposal-for-twitter-tag-channels/, accessed on June 20th, 2023. Twitter adopted hashtags as a feature in 2009 [Scott, 2015], which proved helpful and, later, was also integrated into other social media platforms, such as Facebook [^3] Available at: https://about.fb.com/news/2013/06/public-conversations-on-facebook/. Accessed on June 20th, 2023. in 2013.
Twitter post categorization or grouping leads to virtual community creation. The creation of virtual communities enables discussions on several aspects, and sometimes, online social movements go offline [Smidi & Shahin, 2017]. An example of a hashtag used in social discussion is #mybodymychoice. As pointed out in Menghini et al. [2022], the hashtag #mybodymychoice initially regarded women’s right to abortion and bodily autonomy. However, the authors in Menghini et al. [2022] found that between 2018 and 2021, this hashtag drifted to other topics, such as politics and civil rights.
This paper expands the analysis performed in Menghini et al. [2022] to 2022, considering hashtags co-occurrence and community detection via temporal graphs and textual streaming analysis. In our analysis, we intend to trace back the events that used particular hashtags to understand the hashtags’ occurrences. Therefore, this section presents textual data stream mining and its characteristics, graphs and their notation, and techniques for community detection in graphs.
@@ 2.2 Textual Data Stream Mining
Data streams are “an algorithmic abstraction that allows for real-time analytics” [Gama et al., 2014]. Data arrive sequentially (one by one or in small batches) and quickly in streaming settings. In addition, a data stream may be infinite and ephemeral, meaning that the data distribution may change over time, thus giving rise to a phenomenon named concept drift [Bifet et al., 2018, Gama et al., 2014].
Therefore, to learn from data streams, i.e., data stream mining, some capabilities desired for machine learning methods include [Bifet et al., 2018, Gama et al., 2014]: (a) learning quickly from the data as it arrives, (b) discarding data within a short time after learning from it, (c) performing single-pass processes, and (d) detecting and adapting to concept drifts.
Textual data streams are specializations of traditional data streams [Thuma et al., 2023], i.e., they comprise a sequence of texts arriving continuously at a fast pace. However, processing texts in textual data stream scenarios is more challenging given the natural language processing (NLP)-related activities regarding, for example, vocabulary maintenance, token standardization, and processing texts only once.
Our proposed approach focuses on processing textual data streams while accounting for the constraints mentioned above. The exception regards graph community detection, detailed in Section 2.4, which is performed on a year-basis and, thus, should be considered an offline step.
@@ 2.3 Graph
As stated in Bondy et al. [1976], “many real-world situations can conveniently be described by means of a diagram consisting of a set of points together with lines joining certain pairs of these points”. As examples, these points and lines could represent cities and the roads between them, people and their relationships, or companies and trades between them. The representation based on points and lines is called graph.
A graph $G$ is a collection of finite edges $E$ and finite vertices $V$, generally being represented as $G(V,E)$. Each vertex can be linked to another vertex by an edge. Generally, vertices are visually represented as circles, and the edges are lines between vertices, as depicted in the example graph in Fig. 1. Edges are directed if the direction between vertices matters and are undirected otherwise. Edges may also be coupled with weights to denote additional information about the relationship between connected vertices.
@@
Graphs are relevant to denote entities and their relationships, which are very useful in text analysis. For instance, to find relevant words in sentences, the vertices are the words, and the edges are their relationships [Castillo et al., 2017]. In our proposal, vertices denote hashtags, and edges indicate the co-occurrence between the linked hashtags. The edges are not weighted.
@@ 2.4 Community Detection in Graphs
Community detection methods aim at identifying groups with frequent links, indicating those groups as communities [Daud et al., 2020]. Criminal community detection [Sarvari et al., 2014], cattle trade movement [Cardoso et al., 2018], and disease spread dynamics [Salathé & Jones, 2010] are applications that leverage community detection. Several community detection methods are available in the literature, such as the Louvain method [Blondel et al., 2008], the Leiden algorithm [Traag et al., 2019], and the Girvan-Newman algorithm [Girvan & Newman, 2002].
@@
Similar to clustering, community detection demands the graph partition into densely connected nodes and nodes from a particular community to be far from those of other communities [Blondel et al., 2008]. According to Traag et al. [2019], the Louvain method obtains weakly connected communities. To handle this characteristic, the authors in Traag et al. [2019] added one more phase to the Louvain algorithm to obtain better-connected communities.
@@
The Girvan-Newman algorithm [Girvan & Newman, 2002] for community detection on graphs performs an iterative edge elimination. To consider an edge for elimination, it must be part of the highest number of shortest paths between nodes, i.e., betweenness centrality. After removing these edges, the communities are obtained. More specifically, the algorithm works in four steps [Girvan & Newman, 2002]: (a) for all edges in the graph, compute the betweenness; (b) remove the edge with the highest betweenness; (c) for all edges affected by the edge removal in step (b), re-compute the betweenness; (d) return to step (b) until there is no edge remaining. We resorted to the Girvan-Newman algorithm because it generates consistent communities [Sheng et al., 2020]. However, its trade-off lies in its complexity, which makes it not scalable for large networks [Sheng et al., 2020].
@@
Considering that we are using a textual data stream limited by a window and that the communities are detected in an offline step, community quality is the most important. Therefore, the complexity of the Girvan-Newman algorithm is not a concern in this particular scenario.
2.5 Related Works
As mentioned in Section 2.1, hashtags have been used in social media platforms as a form of post categorization. However, hashtag drift is a recently studied phenomenon. Sometimes, this phenomenon is referred to as hashtag hijacking [VanDam & Tan, 2016] or hashjacking [Darius & Stephany, 2019]. Below, we present some papers that consider hashtag drift and applications.
The authors in Booten [2016] described the hashtag drift based on evolving uses of political hashtags. The authors divided the hashtags into focused and individualistic. Focused hashtags are those which have their use still in the original context, while individualistic hashtags are the hashtags that evolved according to individual’s particular contexts. In this paper, the authors used Tumblr API as the data source. The authors filtered through the API to obtain posts using the most important hashtag in the past years according to a set of web pages. In their experiments, the authors kept nine hashtags, since more than 5,000 posts included at least one of these hashtags. The authors obtained the earliest and latest deciles of posts for each hashtag. The authors also proposed a score to evaluate drifts, considering their observation that states focused hashtags tend to have a small set of frequent co-occurring hashtags, while individualistic hashtags have a “great diversity of infrequent co-occurring hashtags”. The authors conclude by confirming a finding of Bennett & Segerberg [2011], which claimed that “activist movements can survive and thrive by allowing individuals to flexibly re-define the movement’s goals on social media” [Bennett & Segerberg, 2011, Booten, 2016].
Concerned about the challenges of antibiotic resistance, the authors in Mackenzie et al. [2020] evaluated the impact of particular hashtags on disseminating information regarding the overlapping campaigns European Antibiotic Awareness Day and World Antibiotic Awareness Week in 2018. The authors considered a set of 14,400 tweets with 60,222 retweets. In this paper, the authors identified hashtag drift. However, they use a slightly different definition, i.e., “the use and spread of hashtags other than the official hashtags”. One of the findings of the authors stated that an unofficial hashtag #WAAW2018 was a strong predictor of retweet activity for users with less than 1,000 followers.
The authors in Arif et al. [2018] analyzed discourses regarding the #BlackLivesMatter between January and October 2016. This hashtag had its origin in response to the acquittal of George Zimmerman after the shooting death of Trayvon Martin, which happened in 2012. The authors’ filtering resulted in a dataset with 248,719 tweets. When visualizing their dataset, the authors could label them according to the political spectrum. Among their findings, the authors realized a phenomenon close to hashtag drift. For example, some tweets tried to impose new meanings to #BlackLivesMatter, such as “adversarial stance towards law enforcement” [Arif et al., 2018].
In Menghini et al. [2022], the authors analyzed the drift of #MyBodyMyChoice on Twitter, using tweets from January 2018 to December 2021. The authors split the period into three moments, having the Covid-19 pandemic as the reference, i.e., before the pandemic (BP), initial pandemic (IC), and Covid-19 coexistence (CC). The authors noticed a drift/hashjacking related to Covid-19 and vaccination from when the first official Covid-19 case outside China was confirmed. Compared to this case specifically, we evaluated between 2018 and 2022 (inclusive) using the stream paradigm. Herefore, the approach could be directly coupled to a streaming source, such as a social media platform API. In addition, we evaluated the co-occurrences of hashtags to verify drift other than those related to Covid-19.
More recently, the authors in FitzGerald & Graham [2024] evaluated the use of #SaveTheChildren between January 2022 and March 2023 on Twitter (X). The authors assessed that the aforementioned hashtag was used in contexts different from the original. According to the authors, #SaveTheChildren is often associated with charity to aid children and posts related to Covid-19 vaccines. However, the authors also verified that this hashtag is used to spread conspiracy theories related to anti-vaccination and anti-LGBTQIA+ rhetoric. In addition, the authors checked that conspiratorial tweets are more retweeted than non-conspiratorial tweets. In addition, the authors highlight that drifted hashtags can impact legitimate movements, and these drifted/hijacked hashtags can expose more users to conspiracy theories.
As seen, hashtag drift/hijacking, or hashjacking, was identified and has been studied under several aspects, such as political and technical. However, the main difference in our approach is the online construction and maintenance of a hashtag co-occurrence graph with yearly analysis. This characteristic allows the online analysis using a streaming data source. Therefore, with the huge data production and big data scenarios, our approach can be relevant.